# Semantic Refinement with LLMs for Graph Representations

### v1

@@ Abstract
Graph-structured data exhibit substantial heterogeneity in where their predictive signals originate: in some domains, node-level semantics dominate, while in others, structural patterns play a central role. This structure–semantics heterogeneity implies that no graph learning model with a fixed inductive bias can generalize optimally across diverse graph domains. However, most existing methods address this challenge from the model side by incrementally injecting new inductive biases, which remains fundamentally limited given the open-ended diversity of real-world graphs. In this work, we take a data-centric perspective and treat node semantics as a task-adaptive variable. We propose a Data-Adaptive Semantic Refinement framework (DAS) for graph representation learning, which couples a fixed graph neural network (GNN) and a large language model (LLM) in a closed feedback loop. The GNN provides implicit supervisory signals to guide the semantic refinement of LLM, and the refined semantics are fed back to update the same graph learner. We evaluate our approach on both text-rich and text-free graphs. Results show consistent improvements on structure-dominated graphs while remaining competitive on semantics-rich graphs, demonstrating the effectiveness of data-centric semantic adaptation under structure–semantics heterogeneity.
Semantic Refinement with LLMs for Graph Representations
@@ 1 Introduction
Figure 1: Structure–semantics heterogeneity and data-centric adaptation. (a) Real-world graphs vary widely in their reliance on semantic and structural patterns as sources of predictive signal. (b) Model-centric approaches with fixed inductive biases become misaligned when deployed across graphs with different structure–semantics regimes, leading to poor generalization. (c) In contrast, our proposed DAS keeps the graph model fixed and iteratively refines node semantics through prediction-driven feedback, adapting the data representation to the target task.
@@
Graph-structured data (wu2020comprehensive; zhang2020deep) are ubiquitous in the real world, arising in diverse domains such as citation networks, social platforms, molecular interaction systems, and transportation infrastructures. Despite sharing the same graph abstraction, these domains differ fundamentally in where their predictive signals originate. In citation networks, for example, each node represents a scientific document whose topical content and research focus are explicitly encoded in natural language. Here, node-level semantics—captured by titles, abstracts, or full texts—often provide the primary discriminative signal for downstream tasks, while citation links mainly serve as a contextual scaffold that propagates and regularizes semantic information (greenberg2009citation; zhao2015analysis; zhang2019shne). By contrast, in domains such as molecular graphs or transportation networks, semantic attributes are weak or even absent (wu2018moleculenet). Instead, node identity and functionality are determined predominantly by structural roles and global topological patterns, such as motifs, connectivity configurations, and relative positional relationships (chen2020can; zhang2024beyond; wang2025towards). These examples demonstrate that predictive signals in real-world graphs may be dominated by semantics, dominated by structure, or arise from their intricate interplay.
@@
This observation leads to a fundamental and unavoidable consequence: the balance between semantics and structure is inherently domain-dependent, rather than governed by a universal principle. As a result, no graph learning model with a fixed inductive bias can perform optimally across graph domains with drastically different structure–semantics regimes (platonov2023characterizing).
@@
However, translating this observation into a practical learning system remains challenging. For a new graph, the dominant source of predictive signal—whether driven by semantics, structure, or their interaction—is unknown a priori, yet both the model and the data representation must commit to specific inductive biases in advance. Modern GNNs encode fixed architectural preferences once chosen, favoring, for example, locality (velickovic2018graph), long-range dependencies (xu2018how; rampavsek2022recipe), or substructure information (wang2025beyond; wang2025generative). Meanwhile, node representations—whether feature vectors (kipf2017semisupervised), textual embeddings (wang2024gft), or structural descriptors (perozzi2014deepwalk; grover2016node2vec)—are typically constructed in a predefined manner and kept fixed throughout training. As a result, the learning system becomes implicitly specialized to a particular structure–semantics regime. When this specialization is mismatched with the true signal distribution of the target graph, performance degrades systematically, and adaptation in practice is often reduced to empirical model and feature selection rather than a principled mechanism.
@@
To balance semantics and structure, most existing methods approach this problem primarily from the model side. One line of work adapts GNN architectures by redesigning message passing (morris2019weisfeiler; zhang2019heterogeneous; fan2022heterogeneous), incorporating adaptive aggregation (ying2018hierarchical), or injecting positional encodings (murphy2019relational), thereby embedding different inductive biases into the model. Beyond architectural modifications, another line of work introduces external reasoning models (chen2024exploring; wang2023can), most notably large language models (LLMs) (zhao2023survey; ye2025llms4all), which process graph structures and node attributes in textual form. In parallel, other methods rely on auxiliary models (chen2024exploring) to generate additional semantic signals—such as synthetic attributes (he2023harnessing)—that are subsequently consumed by a downstream GNN. Despite their empirical success, these approaches fundamentally rely on incrementally injecting model-level inductive biases, which cannot guarantee universal adaptability across open-ended and structurally diverse graph domains.
@@
In this work, we take a fundamentally different stance on structure-semantics heterogeneity by shifting the adaptation from the model to the data. Instead of continually expanding model-level inductive biases, we treat node semantics as a task-adaptive variable. This shift is motivated by the observation that the balance between structure and semantics is ultimately realized through the input representations consumed by the model, rather than through the architecture alone. As a result, misalignment on new graph domains often arises from fixed node semantics that fail to reflect the graph-specific source of predictive signal.
@@
Building on this perspective, we propose Data-Adaptive Semantic Refinement (DAS), a data-centric, feedback-driven framework for iterative semantic adaptation. Starting from initial node descriptions or structure-derived verbalizations (wang2024tans), we train a fixed GNN for the downstream task and use its predictions as implicit supervision. A large language model then refines node semantics by conditioning on both structural context and model behavior, and the refined descriptions are fed back to update the same graph learner. By iterating this closed loop for a small number of rounds, DAS progressively aligns node semantics with the structure–semantics regime of the target graph without modifying the underlying model. We evaluate DAS on both text-attributed and text-free graphs, where it consistently improves performance on structure-dominated graphs while remaining competitive on semantics-rich graphs.
Figure 2: Overview of the DAS framework. Node descriptions are iteratively refined through a closed loop between a fixed GNN and an LLM. At each iteration, the GNN provides task feedback and a model-conditioned memory retrieves structurally and semantically aligned in-graph exemplars, which guide the LLM to update node semantics before feeding them back to the same GNN.
@@ 2.1 Problem Definition
We consider a graph $G=(\mathcal{V},\mathcal{E})$ with node set $\mathcal{V}$ and edge set $\mathcal{E}$. Each node $v\in\mathcal{V}$ is associated with an initial description $r_{v}$, which is either a natural language text (in text-attributed graphs) or a structure-derived verbalization (in text-free graphs). A subset of nodes $\mathcal{V}_{\mathrm{train}}\subseteq\mathcal{V}$ is labeled with $y_{v}\in\mathcal{Y}$. Given a GNN $g_{\theta}$ for node classification, we treat node semantics as adaptive variables. Our objective is to iteratively refine node descriptions $\{d_{v}\}_{v\in\mathcal{V}}$ from the initial inputs $\{r_{v}\}$, such that the refined semantics better align with both the graph structure and the downstream prediction task. After $T$ refinement steps, the final classifier $g_{\theta}$ trained on $\{d_{v}^{(T)}\}$ is used for evaluation.
@@ 2.2 Overview
DAS is a data-centric, feedback-driven framework that adapts node semantics under structural context and task supervision while keeping the graph model fixed. Instead of modifying model architectures to handle heterogeneous structure–semantics regimes, DAS treats node descriptions as adaptive states that are iteratively refined through a closed loop between a GNN and an LLM.
As illustrated in Figure˜2, DAS operates over $T$ refinement iterations. At iteration $t$, each node $v\in\mathcal{V}$ is associated with a description $d_{v}^{(t)}$ (initialized as $d_{v}^{(0)}=r_{v}$), which is encoded into node features and fed into a fixed GNN $g_{\theta}$ to produce predictions
$\mathbf{p}_{v}^{(t)}=g_{\theta}(\mathbf{x}_{v}^{(t)},G).$ (1)
The current descriptions and predictions are stored in a history buffer $\mathcal{B}^{(t)}$, from which an in-graph support set $\mathcal{S}_{v}^{(t)}$ is retrieved for each node. Conditioned on $d_{v}^{(t)}$ and $\mathcal{S}_{v}^{(t)}$, the LLM refines the node semantics as
$d_{v}^{(t+1)}=\mathcal{M}(d_{v}^{(t)},\mathcal{S}_{v}^{(t)}).$ (2)
The refined descriptions are then fed back to the same GNN, completing one refinement iteration.
Through this feedback-driven loop, DAS progressively aligns node semantics with the structure-semantics regime of the target graph. Unlike prior LLM-enhanced methods that rely on fixed prompts or exemplars (he2023harnessing; chen2024exploring; wang2024tans), DAS enables task-conditioned semantic evolution guided by model behavior.
@@ 2.3 Structure-Aware Initial Node Semantics
We construct initial node descriptions by expressing structural information in natural language, so that both semantic and structural cues can be processed in a unified textual space. Following wang2024tans, for each node $v\in\mathcal{V}$ we compute a small set of structural statistics, including degree, betweenness, closeness, clustering coefficient, and square clustering coefficient (zhang2017degree; saramaki2007generalizations; zhang2008clustering). We present a detailed discussion in Appendix C.
To eliminate scale variation across graphs, each statistic is converted into a percentile rank within the graph. These normalized values are then mapped into a concise structural summary $t_{v}^{\text{struct}}$ via a fixed template, as shown in Appendix I. For text-attributed graphs, we set $d_{v}^{(0)}=[r_{v}\,\|\,t_{v}^{\text{struct}}]$, while for text-free graphs we use $d_{v}^{(0)}=t_{v}^{\text{struct}}$. This design expresses both semantic and structural information in a single textual modality, enabling consistent encoding and subsequent refinement.
@@ 2.4 Model-Conditioned Memory
We maintain a model-conditioned memory to explicitly represent how node semantics, graph structure, and task predictions interact during each refinement stage. At iteration $t$, the memory is denoted as $\mathcal{B}^{(t)}=\{\beta^{(t)}_{v}\}_{v\in\mathcal{V}}$, which stores node-level states induced by the current descriptions $\{d_{v}^{(t)}\}$ under the fixed GNN $g_{\theta}$. For each node $v\in\mathcal{V}$, the memory stores a joint state triple
$\beta^{(t)}_{v}=(d_{v}^{(t)},\,\mathbf{s}_{v},\,\mathbf{p}_{v}^{(t)}),$ (3)
where $d_{v}^{(t)}$ is the current textual description, $\mathbf{s}_{v}$ denotes a structure-oriented embedding encoding the node’s topological role via struc2vec ribeiro2017struc2vec, and $\mathbf{p}_{v}^{(t)}$ is the predictive distribution produced by the GNN. This triple defines a semantic–structural–predictive state for each node under the current representation.
Memory Construction and Update.
At $t=0$, the initial memory $\mathcal{B}^{(0)}$ is constructed from the initial descriptions $\{d_{v}^{(0)}\}$, the fixed structural embeddings $\{\mathbf{s}_{v}\}$, and the GNN predictions obtained using these features. At each subsequent iteration $t>0$, the descriptions are updated to $\{d_{v}^{(t)}\}$, re-encoded as node features, and fed into the same GNN. The resulting predictive states $\{\mathbf{p}_{v}^{(t)}\}$ are then used to overwrite the previous memory entries, yielding an updated memory $\mathcal{B}^{(t)}$. This rolling update scheme ensures that the memory always reflects the latest alignment between node semantics, graph structure, and task-specific behavior.
@@
Given the memory $\mathcal{B}^{(t)}$, the goal of memory retrieval is to identify, for each target node $v$, a small set of in-graph exemplars $\mathcal{S}_{v}^{(t)}$ that are simultaneously semantically relevant, structurally aligned, and reliable under the current classifier. These exemplars serve as task-aware references for subsequent semantic refinement.
To this end, the memory induces a joint semantic–structural similarity space. Let $\mathbf{t}_{v}^{(t)}$ denote the embedding of the current description $d_{v}^{(t)}$ produced by the text encoder (wang2020minilm), and let $\mathbf{s}_{v}$ denote the structural embedding encoding the topological role of node $v$. For any pair of nodes $(v,u)$, we define the semantic similarity
$\text{sim}_{t}(v,u)=\frac{(\mathbf{t}_{v}^{(t)})^{\top}\mathbf{t}_{u}^{(t)}}{\|\mathbf{t}_{v}^{(t)}\|\,\|\mathbf{t}_{u}^{(t)}\|},$ (4)
and the structural similarity
$\text{sim}_{s}(v,u)=\frac{\mathbf{s}_{v}^{\top}\mathbf{s}_{u}}{\|\mathbf{s}_{v}\|\,\|\mathbf{s}_{u}\|}.$ (5)
These two components are combined into a joint similarity score
@@
For each target node $v$, all candidate nodes $u\in\mathcal{V}\setminus\{v\}$ are ranked according to $S(v,u)$. From the top-ranked candidates, we further incorporate model confidence stored in the predictive state $\mathbf{p}_{u}^{(t)}$ to filter unreliable references. Specifically, nodes with low predictive entropy or consistently correct predictions are preferred. The resulting exemplar set $\mathcal{S}_{v}^{(t)}$ thus consists of in-graph references that are not only close to $v$ in the joint semantic–structural space, but also stable with respect to the current task model.
Formally, for each target node $v$, we first rank all candidate nodes $u\in\mathcal{V}\setminus\{v\}$ by the joint similarity score $S(v,u)$. Let $\mathcal{C}_{v}^{(t)}$ denote the top-$K$ candidates under this ranking. We then define a confidence score for each candidate node $u$ based on the predictive distribution $\mathbf{p}_{u}^{(t)}$, for example using the normalized entropy
$H(u)=-\sum_{c\in\mathcal{Y}}p_{u}^{(t)}(c)\,\log p_{u}^{(t)}(c).$ (7)
The final exemplar set is selected as
$\mathcal{S}_{v}^{(t)}=\Big\{\,u\in\mathcal{C}_{v}^{(t)}\;\Big|\;H(u)\leq\tau\,\Big\},$ (8)
where $\tau$ is a confidence threshold. This ensures that selected exemplars are both similar to $v$ in the joint semantic–structural space and reliable under the current classifier.

### v2

@@ Abstract
Graph-structured data exhibit substantial heterogeneity in where their predictive signals originate: in some domains, node-level semantics dominate, while in others, structural patterns play a central role. This structure–semantics heterogeneity implies that no graph learning model with a fixed inductive bias can generalize optimally across diverse graph domains. However, most existing methods address this challenge from the model side by incrementally injecting new inductive biases, which remains fundamentally limited given the open-ended diversity of real-world graphs. In this work, we take a data-centric perspective and treat node semantics as a task-adaptive variable. We propose a Graph-Exemplar-guided Semantic Refinement (GES) framework for graph representation learning which – unlike existing LLM-enhanced methods that generate node descriptions without graph context – leverages structurally and semantically similar nodes from the graph itself to guide semantic refinement. Specifically, a GNN is first trained to produce predictive states, which along with structural and semantic similarity are used to retrieve in-graph exemplars that inform an LLM in refining node descriptions. We evaluate our approach on both text-rich and text-free graphs. Results show consistent improvements on semantics-rich and structure-dominated graphs, demonstrating the effectiveness of data-centric semantic refinement under structure–semantics heterogeneity.
@@ 1 Introduction
Graph-structured data (Wu et al., 2020; Zhang et al., 2020) are ubiquitous in the real world, arising in diverse domains such as citation networks, social platforms, molecular interaction systems, and transportation infrastructures. Despite sharing the same graph abstraction, these domains differ fundamentally in where their predictive signals originate. In citation networks, for example, each node represents a scientific document whose topical content and research focus are explicitly encoded in natural language. Here, node-level semantics—captured by titles, abstracts, or full texts—often provide the primary discriminative signal for downstream tasks, while citation links mainly serve as a contextual scaffold that propagates and regularizes semantic information (Greenberg, 2009; Zhao and Strotmann, 2015; Zhang et al., 2019b). By contrast, in domains such as molecular graphs or transportation networks, semantic attributes are weak or even absent (Wu et al., 2018). Instead, node identity and functionality are determined predominantly by structural roles and global topological patterns, such as motifs, connectivity configurations, and relative positional relationships (Chen et al., 2020; Zhang et al., 2024a; Wang et al., 2025c). These examples demonstrate that predictive signals in real-world graphs may be dominated by semantics, dominated by structure, or arise from their intricate interplay.
@@
This observation leads to a fundamental and unavoidable consequence: the balance between semantics and structure is inherently domain-dependent, rather than governed by a universal principle. As a result, no graph learning model with a fixed inductive bias can perform optimally across graph domains with drastically different structure–semantics regimes (Platonov et al., 2023).
@@
However, translating this observation into a practical learning system remains challenging. For a new graph, the dominant source of predictive signal—whether driven by semantics, structure, or their interaction—is unknown a priori, yet both the model and the data representation must commit to specific inductive biases in advance. Modern GNNs encode fixed architectural preferences once chosen, favoring, for example, locality (Veličković et al., 2018), long-range dependencies (Xu et al., 2019; Rampášek et al., 2022), or substructure information (Wang et al., 2025b; d). Meanwhile, node representations—whether feature vectors (Kipf and Welling, 2017), textual embeddings (Wang et al., 2024b), or structural descriptors (Perozzi et al., 2014; Grover and Leskovec, 2016)—are typically constructed in a predefined manner and kept fixed throughout training. As a result, the learning system becomes implicitly specialized to a particular structure–semantics regime. When this specialization is mismatched with the true signal distribution of the target graph, performance degrades systematically, and adaptation in practice is often reduced to empirical model and feature selection rather than a principled mechanism.
@@
To balance semantics and structure, most existing methods approach this problem primarily from the model side. One line of work adapts GNN architectures by redesigning message passing (Morris et al., 2019; Zhang et al., 2019a; Fan et al., 2022), incorporating adaptive aggregation (Ying et al., 2018), or injecting positional encodings (Murphy et al., 2019), thereby embedding different inductive biases into the model. Beyond architectural modifications, another line of work introduces external reasoning models (Chen et al., 2024b; Wang et al., 2023a), most notably large language models (LLMs) (Zhao et al., 2023b; Ye et al., 2025), which process graph structures and node attributes in textual form. In parallel, other methods rely on auxiliary models (Chen et al., 2024b) to generate additional semantic signals—such as synthetic attributes (He et al., 2023)—that are subsequently consumed by a downstream GNN. Despite their empirical success, these approaches fundamentally rely on incrementally injecting model-level inductive biases, which cannot guarantee universal adaptability across open-ended and structurally diverse graph domains.
Figure 1: Structure–semantics heterogeneity and data-centric adaptation. (a) Real-world graphs vary widely in their reliance on semantic and structural patterns as sources of predictive signal. (b) Model-centric approaches with fixed inductive biases become misaligned when deployed across graphs with different structure–semantics regimes, leading to poor generalization. (c) In contrast, our proposed GES keeps the graph model fixed and refines node semantics through exemplar-guided refinement.
@@
In this work, we take a complementary data-centric perspective on structure-semantics heterogeneity by shifting the adaptation from the model to the data. Instead of continually expanding model-level inductive biases, we treat node semantics as a task-adaptive variable. This shift is motivated by the observation that the balance between structure and semantics is ultimately realized through the input representations consumed by the model, rather than through the architecture alone. As a result, misalignment on new graph domains often arises from fixed node semantics that fail to reflect the graph-specific source of predictive signal.
@@
Building on this perspective, we propose Graph-Exemplar-guided Semantic Refinement (GES), a data-centric framework for exemplar-guided semantic refinement. Starting from initial node descriptions or structure-derived verbalizations (Wang et al., 2025a), we train a GNN for the downstream task and use its predictions as implicit supervision. A large language model then refines node semantics by conditioning on both structural context and model behavior, and the refined descriptions are re-encoded for the final graph learner. Through this single refinement pass, GES aligns node semantics with the structure–semantics regime of the target graph without modifying the underlying model. We evaluate GES on both text-attributed and text-free graphs, where it consistently improves performance on structure-dominated and semantics-rich graphs.
Figure 2: Overview of the GES framework. A GNN is first trained on initial node descriptions to populate a model-conditioned memory, from which structurally and semantically aligned in-graph exemplars are retrieved to guide an LLM in refining node semantics.
@@ 2.1 Problem Definition
We consider a graph $G=(\mathcal{V},\mathcal{E})$ with node set $\mathcal{V}$ and edge set $\mathcal{E}$. Each node $v\in\mathcal{V}$ is associated with an initial description $r_{v}$, which is either a natural language text (in text-attributed graphs) or a structure-derived verbalization (in text-free graphs). A subset of nodes $\mathcal{V}_{\mathrm{train}}\subseteq\mathcal{V}$ is labeled with $y_{v}\in\mathcal{Y}$. Given a GNN $g_{\theta}$ for node classification, we treat node semantics as adaptive variables. Our objective is to refine node descriptions $\{d_{v}\}_{v\in\mathcal{V}}$ from the initial inputs $\{r_{v}\}$, such that the refined semantics better align with both the graph structure and the downstream prediction task. The final classifier $g_{\theta}$ trained on the refined descriptions $\{d_{v}^{\prime}\}$ is used for evaluation.
@@ 2.2 Overview
GES is a data-centric framework that adapts node semantics under structural context and task supervision. Instead of modifying model architectures to handle heterogeneous structure–semantics regimes, GES treats node descriptions as adaptive states that are refined through exemplar-guided LLM rewriting conditioned on GNN feedback.
As illustrated in Figure 2, each node $v\in\mathcal{V}$ starts with an initial description $d_{v}$ (set to $r_{v}$), which is encoded into node features and fed into a GNN $g_{\theta}$ to produce predictions $\mathbf{p}_{v}=g_{\theta}(\mathbf{x}_{v},G)$. The descriptions and predictions are stored in a model-conditioned memory $\mathcal{B}$, from which an in-graph support set $\mathcal{S}_{v}$ is retrieved for each node. Conditioned on $d_{v}$ and $\mathcal{S}_{v}$, the LLM refines the node semantics as $d_{v}^{\prime}=\mathcal{M}(d_{v},\mathcal{S}_{v})$. The refined descriptions are then re-encoded and used to train the final GNN for evaluation. Unlike prior LLM-enhanced methods that rely on fixed prompts or exemplars (He et al., 2023; Chen et al., 2024b; Wang et al., 2025a), GES enables task-conditioned semantic refinement guided by model behavior.
@@ 2.3 Structure-Aware Initial Node Semantics
We construct initial node descriptions by expressing structural information in natural language, so that both semantic and structural cues can be processed in a unified textual space. Following Wang et al. (2025a), for each node $v\in\mathcal{V}$ we compute a small set of structural statistics, including degree, betweenness, closeness, clustering coefficient, and square clustering coefficient (Zhang and Luo, 2017; Saramäki et al., 2007; Zhang et al., 2008). We present a detailed discussion in Appendix C.
To eliminate scale variation across graphs, each statistic is converted into a percentile rank within the graph. These normalized values are then mapped into a concise structural summary $t_{v}^{\text{struct}}$ via a fixed template, as shown in Appendix J. For text-attributed graphs, we set $d_{v}=[r_{v}\,\|\,t_{v}^{\text{struct}}]$, while for text-free graphs we use $d_{v}=t_{v}^{\text{struct}}$. This design expresses both semantic and structural information in a single textual modality, enabling consistent encoding and subsequent refinement.
@@ 2.4 Model-Conditioned Memory
We maintain a model-conditioned memory to explicitly represent how node semantics, graph structure, and task predictions interact. The memory is denoted as $\mathcal{B}=\{\beta_{v}\}_{v\in\mathcal{V}}$, which stores node-level states induced by the initial descriptions $\{d_{v}\}$ under the GNN $g_{\theta}$. For each node $v\in\mathcal{V}$, the memory stores a joint state triple: $\beta_{v}=(d_{v},\,\mathbf{s}_{v},\,\mathbf{p}_{v})$, where $d_{v}$ is the current textual description, $\mathbf{s}_{v}$ denotes a structure-oriented embedding encoding the node’s topological role via struc2vec (Ribeiro et al., 2017), and $\mathbf{p}_{v}$ is the predictive distribution produced by the GNN. This triple defines a semantic–structural–predictive state for each node.
Memory Construction
The memory $\mathcal{B}$ is constructed from the initial descriptions $\{d_{v}\}$, the fixed structural embeddings $\{\mathbf{s}_{v}\}$, and the GNN predictions obtained after training on these features. This construction ensures that the memory reflects the alignment between node semantics, graph structure, and task-specific behavior under the current representation.
@@
Given the memory $\mathcal{B}$, the goal of memory retrieval is to identify, for each target node $v$, a small set of in-graph exemplars $\mathcal{S}_{v}$ that are simultaneously semantically relevant, structurally aligned, and reliable under the current classifier. These exemplars serve as task-aware references for subsequent semantic refinement.
To this end, the memory induces a joint semantic–structural similarity space. Let $\mathbf{t}_{v}$ denote the embedding of the description $d_{v}$ produced by the text encoder (Wang et al., 2020), and let $\mathbf{s}_{v}$ denote the structural embedding encoding the topological role of node $v$. For any pair of nodes $(v,u)$, we define the semantic similarity $\text{sim}_{t}(v,u)=(\mathbf{t}_{v}^{\top}\mathbf{t}_{u})/(\|\mathbf{t}_{v}\|\,\|\mathbf{t}_{u}\|)$ and the structural similarity $\text{sim}_{s}(v,u)=(\mathbf{s}_{v}^{\top}\mathbf{s}_{u})/(\|\mathbf{s}_{v}\|\,\|\mathbf{s}_{u}\|)$. These two components are combined into a joint similarity score
@@
For each target node $v$, all candidate nodes $u\in\mathcal{V}\setminus\{v\}$ are ranked according to $S(v,u)$. From the top-ranked candidates, we further incorporate model confidence stored in the predictive state $\mathbf{p}_{u}$ to filter unreliable references. Specifically, nodes with low predictive entropy are preferred. The resulting exemplar set $\mathcal{S}_{v}$ thus consists of in-graph references that are not only close to $v$ in the joint semantic–structural space, but also stable with respect to the current task model.
Formally, for each target node $v$, we first rank all candidate nodes $u\in\mathcal{V}\setminus\{v\}$ by the joint similarity score $S(v,u)$. Let $\mathcal{C}_{v}$ denote the top-$K$ candidates under this ranking. We then define a confidence score for each candidate node $u$ based on the predictive distribution $\mathbf{p}_{u}$, for example using the normalized entropy $H(u)=-\sum_{c\in\mathcal{Y}}p_{u}(c)\,\log p_{u}(c)$. The final exemplar set is selected as
$\mathcal{S}_{v}=\Big\{\,u\in\mathcal{C}_{v}\;\Big|\;H(u)\leq\tau\,\Big\},$ (2)
where $\tau$ is the entropy threshold that upper-bounds the normalized entropy, retaining only confidently classified exemplars. This ensures that selected exemplars are both similar to $v$ in the joint semantic–structural space and reliable under the current classifier.